OpenStudy (anonymous):

Solve for x: |x| − 8 = −5

OpenStudy (anonymous):

is the answer -3

OpenStudy (anonymous):

-3 and 13

OpenStudy (anonymous):

For aboslute value, you don't know if it is negative or positive

OpenStudy (anonymous):

It can be positive 3, -3. Not 13 because 13 - 8 does not = -5 Absolute value just makes any number positive no matter what

OpenStudy (anonymous):

\[x=\pm3\]

OpenStudy (anonymous):

yeas becaause you have to make the 5 positive and negative

OpenStudy (anonymous):

You add 8 to both sides?

OpenStudy (anonymous):

13 - 8= 5 3 - 8 = -5

OpenStudy (anonymous):

Isolate the x before....

OpenStudy (anonymous):

Haha, you're making this more complicated than you need to :)

OpenStudy (anonymous):

Just isolate your x like in a normal equation and you get l x l = 3

OpenStudy (anonymous):

Your equation is lxl - 8 = -5 You don't change any number except for the number you are taking the absolute value of. So plugging 13 in for x will just equal out to be 5. You can't change the signs of any other number outside of the absolute values parenthesis.